# Constrained Portfolio Analysis in High Dimensions: Tracking Error and Weight Constraints

### v1

@@ 8 Conclusion
A.1 Proof of Theorem 1(i). See that, with $\boldsymbol{\Theta}\mathrel{\mathop{:}}=\boldsymbol{\Sigma}_{y}^{-1}$, $\hat{\boldsymbol{\Theta}}\mathrel{\mathop{:}}=\hat{\boldsymbol{\Sigma}}_{y}^{-1}$, $\boldsymbol{\Theta}$ is symmetric,
@@
Define the scalar $A\mathrel{\mathop{:}}=\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}/p$, and $\hat{A}\mathrel{\mathop{:}}=\boldsymbol{1}_{p}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}/p$, and consider the following right side term in (A.1)
$\left\|\frac{\boldsymbol{\Theta}\boldsymbol{1}_{p}}{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}}-\frac{\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}}{\boldsymbol{1}_{p}^{\prime}\hat{\boldsymbol{\Theta}}\boldsymbol{1}_{p}}\right\|_{1}=\left\|\frac{\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}/p}{\hat{A}}-\frac{\boldsymbol{\Theta}\boldsymbol{1}_{p}/p}{A}\right\|_{1}=\frac{\|(\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}/p)A-(\boldsymbol{\Theta}\boldsymbol{1}_{p}/p)\hat{A}\|_{1}}{|\hat{A}A|}$ (A.2)
@@
By Assumption 7(ii), $A\geq c>0$ for $c>0$ being a positive constant, and by Lemma B.5 of Caner et al. (2023) $|A|\leq C<\infty$, for a positive constant $C>0$. Then by Assumptions 1-7,
@@
where $\hat{\boldsymbol{\Theta}}_{j}\mathrel{\mathop{:}}p\times 1$ is the $j$ th row of $\hat{\boldsymbol{\Theta}}$ shown in column format, and for the rate we use Theorem 2(i) of Caner et al. (2023). Combine all those to have
@@
So by Lemma B.5 of Caner et al. (2023) since $|A|\leq C<\infty$
@@
$\|(\boldsymbol{\Theta}\boldsymbol{1}_{p}/p)(A-\hat{A})\|_{1}\leq|A-\hat{A}|\|\boldsymbol{\Theta}\boldsymbol{1}_{p}/p\|_{1}\leq|A-\hat{A}|\|\boldsymbol{\Theta}\|_{l_{1}},$ (A.9)
by p.345 of Horn and Johnson (2013) for the second inequality. Then by symmetry $\|\boldsymbol{\Theta}\|_{l_{1}}=\|\boldsymbol{\Theta}\|_{l_{\infty}}$, with Assumptions 1-7, and defining $\boldsymbol{L}\mathrel{\mathop{:}}=\boldsymbol{B}[({\bf covf_{t}})^{-1}+\boldsymbol{B}^{\prime}\boldsymbol{\Omega}\boldsymbol{B}]^{-1}\boldsymbol{B}^{\prime}$ and use $\boldsymbol{\Theta}\mathrel{\mathop{:}}=\boldsymbol{\Sigma}_{y}^{-1}$ formula in Section 2
$=$ $O(\sqrt{\bar{s}})+O(\bar{s}r_{T}K^{3/2})=O(\bar{s}r_{T}K^{3/2}),$
where the rates are from (A.96)-(A.97) of Caner et al. (2023). Combine (A.9)(A.10) and (A.4)
$\|(\boldsymbol{\Theta}\boldsymbol{1}_{p}/p)(A-\hat{A})\|_{1}=O_{p}(\bar{s}r_{T}K^{3/2})O_{p}(\bar{s}l_{T})=O_{p}(\bar{s}^{2}l_{T}r_{T}K^{3/2}).$ (A.11)
Combine rates (A.8)(A.11), and clearly (A.11) is slower so the rate in (A.6) is
$\|(\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}/p)A-(\boldsymbol{\Theta}\boldsymbol{1}_{p}/p)\hat{A}\|_{1}=O_{p}(\bar{s}^{2}l_{T}r_{T}K^{3/2}).$ (A.12)
@@
$\frac{\|(\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}/p)A-(\boldsymbol{\Theta}\boldsymbol{1}_{p}/p)\hat{A}\|_{1}}{|\hat{A}A|}=O_{p}(\bar{s}^{2}l_{T}r_{T}K^{3/2})=o_{p}(1),$ (A.13)
by Assumption 7(i), $\bar{s}^{2}l_{T}r_{T}K^{3/2}=o(1)$. Define scalar $F\mathrel{\mathop{:}}=\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}/p$, and $\hat{F}\mathrel{\mathop{:}}=\boldsymbol{1}_{p}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\hat{\boldsymbol{\mu}}/p$ on right side term in (A.1)
@@
where we use Holder’s inequality for the first inequality, and p.345 of Horn and Johnson (2013) for the second inequality. Note that $\boldsymbol{\Theta}$ is symmetric
$\left|\frac{\boldsymbol{1}_{p}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\hat{\boldsymbol{\mu}}}{p}-\frac{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}^{\prime}\boldsymbol{\mu}}{p}\right|$$\leq$ $\frac{1}{p}\left|\boldsymbol{1}_{p}^{\prime}(\hat{\boldsymbol{\Theta}}-\boldsymbol{\Theta})^{\prime}(\hat{\boldsymbol{\mu}}-\boldsymbol{\mu})\right|+\frac{1}{p}\left|\boldsymbol{1}_{p}(\hat{\boldsymbol{\Theta}}-\boldsymbol{\Theta})^{\prime}\boldsymbol{\mu}\right|+\frac{1}{p}\left|\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}^{\prime}(\hat{\boldsymbol{\mu}}-\boldsymbol{\mu})\right|$
@@
where we use (A.16) and the same technical analysis in (A.4), and the rates are from (A.7) here and then Theorem 2(ii) of Caner et al. (2023) for the first term on the right side, and for the second term rates are by (A.7) here and then (B.8) of Caner et al. (2023), and for the third term, rates are by (A.10) here and $\|\boldsymbol{\Theta}^{\prime}\|_{l_{1}}=\|\boldsymbol{\Theta}\|_{l_{\infty}}$, and Theorem 2(ii) of Caner et al. (2023). Second rate is the slowest by $l_{T}$ definition in (2).
By Lemma B.5 of Caner et al. (2023)
$|F|=O(K^{1/2}),$
and by the condition $|F|\geq c>0$, we have by Assumption 7
@@
$=$ $\|\boldsymbol{\Theta}\|_{l_{\infty}}\|\hat{\boldsymbol{\mu}}-\boldsymbol{\mu}\|_{\infty}=O(\bar{s}r_{T}K^{3/2})O_{p}(max(K\sqrt{\frac{lnT}{T}},\sqrt{\frac{lnp}{T}})))$
where we use the same analysis as in (A.20) for the inequality, and the rates are by (A.10) here and Theorem 2(ii)of Caner et al. (2023). Then (A.20) is slower as a rate than in (A.21) due to $l_{T}$ definition in(2). So with $|F|=O(\sqrt{K})$ by Lemma B.5 of Caner et al. (2023), in (A.18) the first right side term is
@@
$\leq$ $\|\boldsymbol{\Theta}\|_{l_{1}}\|\boldsymbol{\mu}\|_{\infty}|\hat{F}-F|$
$=$ $\|\boldsymbol{\Theta}\|_{l_{\infty}}\|\boldsymbol{\mu}\|_{\infty}|\hat{F}-F|$
$=$ $O_{p}(\bar{s}r_{T}K^{3/2})O(K)O_{p}(K\bar{s}l_{T}),$
where we use p.345 of Horn and Johnson (2013) and $\|\mu\|_{1}\leq p\|\mu\|_{\infty}$ for the first inequality, and the rates are by (A.10) here, and then by (B.8) of Caner et al. (2023) which is $\|\mu\|_{\infty}=O(K)$ and (A.15). Clearly the rate in (A.23) is slower than the one in (A.22), so (A.18)
$\|\frac{\hat{\boldsymbol{\Theta}}^{\prime}\hat{\boldsymbol{\mu}}}{p}F-\frac{\boldsymbol{\Theta}\boldsymbol{\mu}}{p}\hat{F}\|_{1}=O_{p}(\bar{s}^{2}l_{T}r_{T}K^{7/2}).$ (A.24)
@@
$\left\|\left(\frac{\hat{\boldsymbol{\Theta}}^{\prime}\hat{\boldsymbol{\mu}}}{1_{p}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\hat{\boldsymbol{\mu}}}-\frac{\boldsymbol{\Theta}\boldsymbol{\mu}}{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}}\right)\right\|_{1}=O_{p}(\bar{s}^{2}l_{T}r_{T}K^{7/2})=o_{p}(1),$ (A.25)
by Assumption 7(i) to have the last equality. Use (A.13)(A.24) in the weights definition in (A.1)
$\|\hat{\boldsymbol{w}}_{d}-\boldsymbol{w}_{d}^{*}\|_{1}=O_{p}(\bar{s}^{2}l_{T}r_{T}K^{7/2})=o_{p}(1)$
@@
$\hat{\boldsymbol{w}}^{\prime}\boldsymbol{\Sigma}_{y}\hat{\boldsymbol{w}}-\boldsymbol{w}^{*^{\prime}}\boldsymbol{\Sigma}_{y}\boldsymbol{w}^{*}=(\hat{\boldsymbol{w}}_{d}^{\prime}\boldsymbol{\Sigma}_{y}\hat{\boldsymbol{w}}_{d}-\boldsymbol{w}_{d}^{*^{\prime}}\boldsymbol{\Sigma}_{y}\boldsymbol{w}_{d}^{*})+2\boldsymbol{m}^{\prime}\boldsymbol{\Sigma}_{y}(\hat{\boldsymbol{w}}_{d}-\boldsymbol{w}_{d}).$ (A.26)
@@
$\|\boldsymbol{\Sigma}_{y}\|_{\infty}$$\leq$ $\|\boldsymbol{B}{\bf covf_{t}}\boldsymbol{B}^{\prime}\|_{\infty}+\|\boldsymbol{\Sigma}_{u}\|_{\infty}$ (A.28)
$\leq$ $K^{2}\|\boldsymbol{B}\|_{\infty}^{2}\|{\bf covf_{t}}\|_{\infty}+\|\boldsymbol{\Sigma}_{u}\|_{\infty}=O(K^{2}),$
where we obtain the second inequality using Lemma A.1(ii) of Caner et al. (2023) which is $\|\boldsymbol{A}_{1}\boldsymbol{B}_{1}\boldsymbol{A}_{1}^{\prime}\|_{\infty}\leq K^{2}\|\boldsymbol{A}_{1}\|_{\infty}\|\boldsymbol{B}_{1}\|_{\infty}$, where $\boldsymbol{A}_{1}\mathrel{\mathop{:}}p\times K$ matrix, $\boldsymbol{B}_{1}\mathrel{\mathop{:}}K\times K$ matrix, and the rates are by Assumption 1, 4, 5(i).
@@
First we consider the denominators in (A.31). By Assumption 7(ii) and $|\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}|/p\geq c>0$ in the statement of Theorem 1, denominators are bounded away from zero. For the numerators in (A.31), now see that $\boldsymbol{\Theta}$ is symmetric, and by (A.10), with p.345 of Horn and Johnson (2013), and by (B.8) of Caner et al. (2023) since $\|\boldsymbol{\mu}\|_{\infty}=O(K)$
$\|\boldsymbol{\Theta}\boldsymbol{\mu}/p\|_{1}$$\leq$ $\frac{1}{p}\|\boldsymbol{\Theta}\|_{l_{1}}\|\boldsymbol{\mu}\|_{1}$ (A.32)
$\leq$ $\|\boldsymbol{\Theta}\|_{l_{\infty}}\|\boldsymbol{\mu}\|_{\infty}=O(\bar{s}r_{T}K^{3/2})O(K)=O(\bar{s}r_{T}K^{5/2}).$
@@
$\|\boldsymbol{\Theta}\boldsymbol{1}_{p}/p\|_{1}=O(\bar{s}r_{T}K^{3/2}).$ (A.33)
@@
$|\boldsymbol{w}_{d}^{*^{\prime}}\boldsymbol{\Sigma}_{y}(\hat{\boldsymbol{w}}_{d}-\boldsymbol{w}_{d}^{*})|=O_{p}(r_{w1}\bar{s}r_{T}K^{9/2})=o_{p}(1),$ (A.36)
@@
$|\hat{\boldsymbol{w}}_{d}^{\prime}\boldsymbol{\Sigma}_{y}\hat{\boldsymbol{w}}_{d}-\boldsymbol{w}_{d}^{*^{\prime}}\boldsymbol{\Sigma}_{y}\boldsymbol{w}_{d}^{*}|=O_{p}(r_{w1}\bar{s}r_{T}K^{9/2})=o_{p}(1),$ (A.37)
by $r_{w1}$ rate definition and Assumption 7(i). Since we assume $\|\boldsymbol{m}\|_{1}=O(\|w_{d}^{*}\|_{1})$, we have
$|\boldsymbol{m}^{\prime}\boldsymbol{\Sigma}_{y}(\hat{\boldsymbol{w}}_{d}-\boldsymbol{w}_{d}^{*})|=O_{p}(r_{w1}\bar{s}r_{T}K^{9/2})=o_{p}(1),$ (A.38)
where we use exactly the same analysis in (A.36). Next combine (A.37)(A.38) in (A.26), and Assumption 7(iii) to have the desired result.Q.E.D
@@
$\left|\left[\frac{\widehat{SR}}{SR}\right]^{2}-1\right|$$\leq$ $\left|\frac{(\hat{\boldsymbol{w}}^{\prime}\boldsymbol{\mu})^{2}-(\boldsymbol{w}^{*^{\prime}}\boldsymbol{\mu})^{2}}{(\boldsymbol{w}^{*^{\prime}}\boldsymbol{\mu})^{2}}\right|\left|\frac{\boldsymbol{w}^{*^{\prime}}\boldsymbol{\Sigma}_{y}\boldsymbol{w}^{*}-\hat{\boldsymbol{w}}^{\prime}\boldsymbol{\Sigma}_{y}\hat{\boldsymbol{w}}}{\hat{\boldsymbol{w}}^{\prime}\boldsymbol{\Sigma}_{y}\hat{\boldsymbol{w}}}\right|$ (A.39)
$+$ $\left|\frac{(\hat{\boldsymbol{w}}^{\prime}\boldsymbol{\mu})^{2}-(\boldsymbol{w}^{*^{\prime}}\boldsymbol{\mu})^{2}}{(\boldsymbol{w}^{*^{\prime}}\boldsymbol{\mu})^{2}}\right|+\left|\frac{\boldsymbol{w}^{*^{\prime}}\boldsymbol{\Sigma}_{y}\boldsymbol{w}^{*}-\hat{\boldsymbol{w}}^{\prime}\boldsymbol{\Sigma}_{y}\hat{\boldsymbol{w}}}{\hat{\boldsymbol{w}}^{\prime}\boldsymbol{\Sigma}_{y}\hat{\boldsymbol{w}}}\right|.$
@@
$=$ $O_{p}(r_{w1}K)+O(\bar{s}r_{T}K^{7/2})=o_{p}(1)+O(\bar{s}r_{T}K^{7/2})=O_{p}(\bar{s}r_{T}K^{7/2}),$
where for the first inequality we use Holder’s, and then the second inequality is by triangle inequality, for the rates we use Theorem 1(i) here and (B.8) of Caner et al. (2023) in which $\|\boldsymbol{\mu}\|_{\infty}=O(K)$, and $\boldsymbol{w}^{*}\mathrel{\mathop{:}}=\boldsymbol{w}_{d}^{*}+\boldsymbol{m}$, (A.34), Assumption 7(iv), and the last rate is by Assumption 7(i) with $r_{w1}=o(1)$ definition.
@@
$\frac{|(\hat{\boldsymbol{w}}^{\prime}\boldsymbol{\mu})^{2}-(\boldsymbol{w}^{*^{\prime}}\boldsymbol{\mu})^{2}|}{(\boldsymbol{w}^{*^{\prime}}\boldsymbol{\mu})^{2}}=O_{p}(r_{w1}K)O_{p}(\bar{s}r_{T}K^{7/2})=O_{p}(\bar{s}^{3}r_{T}^{2}l_{T}K^{8})=o_{p}(1),$ (A.43)
@@
$\left|\frac{\boldsymbol{w}^{*^{\prime}}\boldsymbol{\Sigma}_{y}\boldsymbol{w}^{*}-\hat{\boldsymbol{w}}^{\prime}\boldsymbol{\Sigma}_{y}\hat{\boldsymbol{w}}}{\hat{\boldsymbol{w}}^{\prime}\boldsymbol{\Sigma}_{y}\hat{\boldsymbol{w}}}\right|$$\leq$ $\left|\frac{\boldsymbol{w}^{*^{\prime}}\boldsymbol{\Sigma}_{y}\boldsymbol{w}^{*}-\hat{\boldsymbol{w}}^{\prime}\boldsymbol{\Sigma}_{y}\hat{\boldsymbol{w}}}{\boldsymbol{w}^{*^{\prime}}\boldsymbol{\Sigma}_{y}\boldsymbol{w}^{*}-|\hat{\boldsymbol{w}}^{\prime}\boldsymbol{\Sigma}_{y}\hat{\boldsymbol{w}}-\boldsymbol{w}^{*^{\prime}}\boldsymbol{\Sigma}_{y}\boldsymbol{w}^{*}|}\right|$ (A.44)
$=$ $O_{p}(r_{w1}\bar{s}r_{T}K^{9/2})=O_{p}(\bar{s}^{3}r_{T}^{2}l_{T}K^{8}),$
@@
A.2 Relation Between Tracking Error and Risk Aversion

### v2

@@ A.2 Proof of Theorem 1
Proof of Theorem 1(i). See that, with $\boldsymbol{\Theta}:=\boldsymbol{\Sigma}_{y}^{-1}$, $\hat{\boldsymbol{\Theta}}:=\hat{\boldsymbol{\Sigma}}_{y}^{-1}$, $\boldsymbol{\Theta}$ is symmetric,
@@
Define the scalar $A:=\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}/p$, and $\hat{A}:=\boldsymbol{1}_{p}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}/p$, and consider the following right side term in (A.2)
$\left\|\frac{\boldsymbol{\Theta}\boldsymbol{1}_{p}}{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}}-\frac{\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}}{\boldsymbol{1}_{p}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}}\right\|_{1}=\left\|\frac{\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}/p}{\hat{A}}-\frac{\boldsymbol{\Theta}\boldsymbol{1}_{p}/p}{A}\right\|_{1}=\frac{\|(\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}/p)A-(\boldsymbol{\Theta}\boldsymbol{1}_{p}/p)\hat{A}\|_{1}}{|\hat{A}A|}$ (A.3)
@@
By Assumption 7(ii), $A\geq c>0$ for $c>0$ being a positive constant, and by Lemma B.5 of Caner et al. (2023) $A\leq C<\infty$, for a positive constant $C>0$. Then by Assumptions 1-7,
@@
where $\hat{\boldsymbol{\Theta}}_{j}:p\times 1$ is the $j$ th row of $\hat{\boldsymbol{\Theta}}$ shown in column format, and for the rate we use Theorem 2(i) of Caner et al. (2023). Combine all those to have
@@
So by Lemma B.5 of Caner et al. (2023) since $A\leq C<\infty$
@@
$A^|,\|(\boldsymbol{\Theta}\boldsymbol{1}_{p}/p)(A-\hat{A})\|_{1}\leq|A-\hat{A}|\|\boldsymbol{\Theta}\boldsymbol{1}_{p}/p\|_{1}\leq C$|A-$\hat${A}|, (A.10)
by p.345 of Horn and Johnson (2013) for the second inequality, and we use
$p^{-1}\|\boldsymbol{\Theta}\boldsymbol{1}_{p}\|_{1}\leq p^{-1/2}\|\boldsymbol{\Theta}\boldsymbol{1}_{p}\|_{2}\leq p^{-1/2}\|\boldsymbol{1}_{p}\|_{2}Eigmax(\boldsymbol{\Theta})\leq C,$ (A.11)
where we use $\boldsymbol{\Theta}$ symmetric and Exercise 7.53b of Abadir and Magnus (2005) for the second inequality, and third inequality is by Lemma B.5 of Caner et al. (2023).
Combine (A.10) (A.5)
$\|(\boldsymbol{\Theta}\boldsymbol{1}_{p}/p)(A-\hat{A})\|_{1}=O(1)O_{p}(\bar{s}l_{T})=O_{p}(\bar{s}l_{T}).$ (A.12)
Combine rates (A.9)(A.12), so the rate in (A.7) is
$\|(\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}/p)A-(\boldsymbol{\Theta}\boldsymbol{1}_{p}/p)\hat{A}\|_{1}=O_{p}(\bar{s}l_{T}).$ (A.13)
@@
$\frac{\|(\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}/p)A-(\boldsymbol{\Theta}\boldsymbol{1}_{p}/p)\hat{A}\|_{1}}{|\hat{A}A|}=O_{p}(\bar{s}l_{T})=o_{p}(1),$ (A.14)
by Assumption 7(i), $\bar{s}l_{T}=o(1)$. Define scalar $F:=\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}/p$, and $\hat{F}:=\boldsymbol{1}_{p}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\hat{\boldsymbol{\mu}}/p$ on right side term in (A.2)
@@
where we use Holder’s inequality for the first inequality, and p.345 of Horn and Johnson (2013) for the second inequality. Before the next proof we have the following result. By symmetry $\|\boldsymbol{\Theta}\|_{l_{1}}=\|\boldsymbol{\Theta}\|_{l_{\infty}}$, with Assumptions 1-7, and defining $\boldsymbol{L}:=\boldsymbol{B}[({\bf\Sigma_{f}})^{-1}+\boldsymbol{B}^{\prime}\boldsymbol{\Omega}\boldsymbol{B}]^{-1}\boldsymbol{B}^{\prime}$ and use $\boldsymbol{\Theta}:=\boldsymbol{\Sigma}_{y}^{-1}$ formula in Section 2
$=$ $O(\sqrt{\bar{s}})+O(\bar{s}r_{T}K^{3/2})=O(\bar{s}r_{T}K^{3/2}),$
where the rates are from (A.96)-(A.97) of Caner et al. (2023). Note that $\boldsymbol{\Theta}$ is symmetric, and adding and subtracting
$\left|\frac{\boldsymbol{1}_{p}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\hat{\boldsymbol{\mu}}}{p}-\frac{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}}{p}\right|$ $\leq$ $\frac{1}{p}\left|\boldsymbol{1}_{p}^{\prime}(\hat{\boldsymbol{\Theta}}-\boldsymbol{\Theta})^{\prime}(\hat{\boldsymbol{\mu}}-\boldsymbol{\mu})\right|+\frac{1}{p}\left|\boldsymbol{1}_{p}(\hat{\boldsymbol{\Theta}}-\boldsymbol{\Theta})^{\prime}\boldsymbol{\mu}\right|+\frac{1}{p}\left|\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}(\hat{\boldsymbol{\mu}}-\boldsymbol{\mu})\right|$ (A.19)
@@
where we use (A.17) and the same technical analysis in (A.5), and the rates are by using Holder’s inequality and from (A.8) here and then Theorem 2(ii) of Caner et al. (2023) for the first term on the right side, and for the second term rates are by (A.8) here and then (B.8) of Caner et al. (2023) which is $\|\boldsymbol{\mu}\|_{\infty}=O(K)$ , and for the third term, rates are by (A.18) here and $\|\boldsymbol{\Theta}^{\prime}\|_{l_{1}}=\|\boldsymbol{\Theta}\|_{l_{\infty}}$, and Theorem 2(ii) of Caner et al. (2023). Second rate is the slowest by $l_{T}$ definition in (2) in main text.
By Cauchy Schwartz inequality
$|F|\leq\left|\frac{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}}{p}\right|^{1/2}\left|\frac{\boldsymbol{\mu}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}}{p}\right|^{1/2}=O(K^{1/2}),$ (A.20)
where the last inequality is by Lemma B.5 of Caner et al. (2023). Then by the condition $|F|\geq c>0$, we have by Assumption 7 and by (A.16)
@@
$=$ $\|\boldsymbol{\Theta}\|_{l_{\infty}}\|\hat{\boldsymbol{\mu}}-\boldsymbol{\mu}\|_{\infty}$
$=$ $O(\bar{s}r_{T}K^{3/2})O_{p}(max(K\sqrt{\frac{lnT}{T}},\sqrt{\frac{lnp}{T}})))$
where we use the same analysis as in (A.24) for the inequality, and the rates are by (A.18) here and Theorem 2(ii) of Caner et al. (2023). Then (A.24) is slower as a rate than in (A.25) due to $l_{T}$ definition in (2). So with $|F|=O(\sqrt{K})$, in (A.22) the first right side term is
@@
$\leq$ $\frac{1}{p^{1/2}}\|\boldsymbol{\Theta}\boldsymbol{\mu}\|_{2}|\hat{F}-F|\leq[Eigmax(\boldsymbol{\Theta})]^{1/2}\left|\frac{\boldsymbol{\mu}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}}{p}\right|^{1/2}|\hat{F}-F|$
$=$ $O(K^{1/2})O_{p}(K\bar{s}l_{T})=O_{p}(\bar{s}l_{T}K^{3/2}),$
where we use Exercise 7.53b of Abadir and Magnus (2005) and and Lemma B.5 of Caner et al. (2023). Clearly the rate in (A.22)
$\|\frac{\hat{\boldsymbol{\Theta}}^{\prime}\hat{\boldsymbol{\mu}}}{p}F-\frac{\boldsymbol{\Theta}\boldsymbol{\mu}}{p}\hat{F}\|_{1}=O_{p}(\bar{s}l_{T}K^{3/2}).$ (A.28)
@@
$\left\|\left(\frac{\hat{\boldsymbol{\Theta}}^{\prime}\hat{\boldsymbol{\mu}}}{1_{p}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\hat{\boldsymbol{\mu}}}-\frac{\boldsymbol{\Theta}\boldsymbol{\mu}}{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}}\right)\right\|_{1}=O_{p}(\bar{s}l_{T}K^{3/2})=o_{p}(1),$ (A.29)
by Assumption 7(i) to have the last equality. Use (A.14)(A.29) in the weights definition in (A.2), with $r_{w1}:=\bar{s}l_{T}K^{3/2}$,
$\|\hat{\boldsymbol{w}}_{d}-\boldsymbol{w}_{d}^{*}\|_{1}=O_{p}(\bar{s}l_{T}K^{3/2})=O_{p}(r_{w1})=o_{p}(1)$
@@
$\hat{\boldsymbol{w}}^{\prime}\boldsymbol{\Sigma}_{y}\hat{\boldsymbol{w}}-\boldsymbol{w}^{*^{\prime}}\boldsymbol{\Sigma}_{y}\boldsymbol{w}^{*}=(\hat{\boldsymbol{w}}_{d}^{\prime}\boldsymbol{\Sigma}_{y}\hat{\boldsymbol{w}}_{d}-\boldsymbol{w}_{d}^{*^{\prime}}\boldsymbol{\Sigma}_{y}\boldsymbol{w}_{d}^{*})+2\boldsymbol{m}^{\prime}\boldsymbol{\Sigma}_{y}(\hat{\boldsymbol{w}}_{d}-\boldsymbol{w}_{d}^{*}).$ (A.30)
@@
$\|\boldsymbol{\Sigma}_{y}\|_{\infty}$ $\leq$ $\|\boldsymbol{B}{\bf\Sigma}_{f}\boldsymbol{B}^{\prime}\|_{\infty}+\|\boldsymbol{\Sigma}_{u}\|_{\infty}$ (A.32)
$\leq$ $K^{2}\|\boldsymbol{B}\|_{\infty}^{2}\|{\bf\Sigma}_{f}\|_{\infty}+\|\boldsymbol{\Sigma}_{u}\|_{\infty}=O(K^{2}),$
where we obtain the second inequality using Lemma A.1(ii) of Caner et al. (2023) which is $\|\boldsymbol{A}_{1}\boldsymbol{B}_{1}\boldsymbol{A}_{1}^{\prime}\|_{\infty}\leq K^{2}\|\boldsymbol{A}_{1}\|_{\infty}^{2}\|\boldsymbol{B}_{1}\|_{\infty}$, where $\boldsymbol{A}_{1}:p\times K$ matrix, $\boldsymbol{B}_{1}:K\times K$ matrix, and the rates are by Assumption 1, 4, 5(i).
@@
First we consider the denominators in (A.35). By Assumption 7(ii) and $|\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}|/p\geq c>0$ in the statement of Theorem 1, denominators are bounded away from zero. For the numerators in (A.35), as in (A.27)
$\|\boldsymbol{\Theta}\boldsymbol{\mu}/p\|_{1}=O(K^{1/2}).$ (A.36)
@@
$\|\boldsymbol{\Theta}\boldsymbol{1}_{p}/p\|_{1}=O(1).$ (A.37)
@@
$|\boldsymbol{w}_{d}^{*^{\prime}}\boldsymbol{\Sigma}_{y}(\hat{\boldsymbol{w}}_{d}-\boldsymbol{w}_{d}^{*})|=O_{p}(r_{w1}K^{5/2})=o_{p}(1),$ (A.40)
@@
$|\hat{\boldsymbol{w}}_{d}^{\prime}\boldsymbol{\Sigma}_{y}\hat{\boldsymbol{w}}_{d}-\boldsymbol{w}_{d}^{*^{\prime}}\boldsymbol{\Sigma}_{y}\boldsymbol{w}_{d}^{*}|=O_{p}(r_{w1}K^{5/2})=o_{p}(1),$ (A.41)
by $r_{w1}$ rate definition and Assumption 7(i). Since we assume $\|\boldsymbol{m}\|_{1}=O(\|\boldsymbol{w}_{d}^{*}\|_{1})$, we have
$|\boldsymbol{m}^{\prime}\boldsymbol{\Sigma}_{y}(\hat{\boldsymbol{w}}_{d}-\boldsymbol{w}_{d}^{*})|=O_{p}(r_{w1}K^{5/2})=o_{p}(1),$ (A.42)
where we use exactly the same analysis in (A.40). Next combine (A.41)(A.42)(A.33) in (A.30), and Assumption 7(iii) to have the desired result.Q.E.D
@@
[…]
@@
$=$ $O_{p}(r_{w1}K)+O(K^{3/2})=o_{p}(1)+O(K^{3/2})$
$=$ $O_{p}(K^{3/2}),$
where for the first inequality we use Holder’s, and then the second inequality is by triangle inequality, for the rates we use Theorem 1(i) here and (B.8) of Caner et al. (2023) in which $\|\boldsymbol{\mu}\|_{\infty}=O(K)$, and $\boldsymbol{w}^{*}:=\boldsymbol{w}_{d}^{*}+\boldsymbol{m}$, (A.38), Assumption 7(iv), and the last rate is by Assumption 7(i) with $r_{w1}=o(1)$ definition, and $\bar{s},K$ as nondecreasing functions of $T$.
@@
$\frac{|(\hat{\boldsymbol{w}}^{\prime}\boldsymbol{\mu})^{2}-(\boldsymbol{w}^{*^{\prime}}\boldsymbol{\mu})^{2}|}{(\boldsymbol{w}^{*^{\prime}}\boldsymbol{\mu})^{2}}=O_{p}(r_{w1}K)O_{p}(K^{3/2})=O_{p}(\bar{s}l_{T}K^{4})=o_{p}(1),$ (A.47)
@@
[…]
$=$ $O_{p}(r_{w1}K^{5/2})=O_{p}(\bar{s}l_{T}K^{4}),$
@@
A.3 Relation Between Tracking Error and Risk Aversion